# ZMM-TTS: Zero-shot Multilingual and Multispeaker Speech Synthesis Conditioned on Self-supervised Discrete Speech Representations

### v1

@@ III METHOD
Figure 1: Overview of ZMM-TTS. The modules txt2vec and vec2wav are trained independently.
Our proposed model is a multilingual and multispeaker synthesis system that uses self-supervised discrete speech representations from a pre-trained multilingual wav2vec 2.0 (XLSR-53) model. Quantization in XLSR-53 is based on product quantization by choosing quantized representations from codebooks $\bm{C}\in\mathbb{R}^{G\times M\times D}$ with $M=320$ entries each, $G=2$ are two separate codebooks in XLSR-53 and each quantization representation has $D=384$ dimsensions.
Given a dataset $\mathcal{S}=\{x_{i},y_{i}\}$, where $y$ is an audio sample and $x=\{x_{0},x_{1},...,x_{T_{m}}\}$ is its corresponding text sequences, we use a pre-trained XLSR-53 model to encode each audio sample into discrete code index sequences $V=\{{v_{1}^{1:G},v_{2}^{1:G},...,v_{T_{c}}^{1:G}\}}$ and the corresponding discrete representation is $R=\{r_{1}^{1:G},r_{2}^{1:G},...,r_{T_{c}}^{1:G}\}$. Obviously, there is a one-to-one correspondence between $V$ and $R$, and once we obtain $V$, we can obtain $R$ by looking up codebooks $\bm{C}$ by $r=C[:,v,:]$.
Fig.1 depicts the architectures in which our proposed model can be implemented. It consists of two steps, as shown in the following formula. First, the discrete code index sequences $V$ are predicted from the text $x$ and the codebook of XLSR-53 $\bm{C}$ is looked up to obtain the discrete representations $R$, and then the waveform is predicted from the discrete representations $R$.
@@
where $S$ denotes the speaker representation and $L$ represents the language identity. The txt2vec module is similar to the acoustic model in prior speech synthesis approaches, so we constructed it on the basis of the popular acoustic model FastSpeech. Vec2wav is more similar to a vocoder. To obtain a broader understanding of the potential complementarity and differences between SSL representations and Mel features, we proposed two distinct approaches to implement vec2wav. The first is to convert discrete representations $R$ into Mel spectrograms and then convert them through an additional vocoder, and the second is to directly convert $R$ into audio. Both methods make use of the up-sampling operation based on transposed convolution in HiFi-GAN [51], as in [44]. It should be noted that although there are $G$ different codebooks, in the vec2wav model, the input $R$ is represented by flattened representations from $G$ codebooks together. In the following two subsections, we will provide an explanation of the two components, respectively.
@@ III-A Predicting Discrete Code Index from Text
Figure 2: The architecture of our proposed text2vec model.
The txt2vec model is based on FastSpeech [31], a non-autoregressive (NAR) acoustic model with explicit duration modeling. In our work, we use a learnable aligner [52] instead of relying on the predicted phonemes durations of an autoregressive teacher TTS model. This aligner is more robust for long sequences. The decoder part is the same as that of FastSpeech, which is a Feed-Forward Transformer (FFT). As for the encoder part, depending on the different types of input representations, it can be either FFT-like or BERT-like. Fig. 2 shows the structure of txt2vec.
@@ III-A1 Input representations and text encoder
We try different input representations including characters, IPA, and pre-trained phoneme representations for multilingual TTS.
@@
III-A2 Multi-speaker and multilingual control
@@ III-A3 Loss Function
In this text2vec stage, the text encoder operates on the input tokens $x$ and produces hidden states, which, combined with language and speaker embeddings, are used to predict the duration $d$ by the duration predictor. The upsampled encoder outputs, together with the language and speaker embeddings, are used as the decoder input. The decoder then generates discrete values for both $V$ and $R$.
@@
$\begin{split}\hat{r}_{t}^{g}=\sum_{i=1}^{V}p(v_{t}^{g}=i|h_{t}^{g})C_{g,i}\end{split}$ (4)
and the regression L2 loss $\mathcal{L}_{MSE}$ between the predicted and targeted $R$ could be represented as
$\mathcal{L}_{mse}=\frac{1}{T}\sum_{t=1}^{T}\frac{1}{G}\sum_{g=1}^{G}{\left\|\hat{r}_{t}^{g},r_{t}^{g}\right\|}^{2}_{2}$ (5)
@@
Note that since $p(v_{t}^{g})$ is a one-hot vector, the $r_{t}^{g}$ is only one entry in the codebook instead of being multiplied by softmax weights as in Eq. (5). Therefore, in inference, we directly select the maximum probability index in the $P(i|h_{t}^{g})$, and look up the corresponding $(\hat{r}_{t}^{g})$ in the code $C^{G*M*D}$ as the input of the next-stage model:
@@
III-B Predicting waveform from discrete representations
Although most methods convert features directly into audio [44], spectral features may still have several advantages in some low-resource tasks [53]. To obtain a broader understanding of the potential complementarity and differences between SSL representations and Mel features, we adopted two different ways to implement the vec2wav model: with and without an independent Mel-based vocoder as in the following formula,
@@ III-B Predicting waveform from discrete representations
This vec2wav model consists of four parts: up-sampling, down-sampling, decoder, and pre-trained Mel-based vocoder as shown in Fig. 3. To learn Mel spectrograms from $R$, the duration model is no longer necessary. Although the sequences $R$ and $Mel$ have different “sampling rates,” they can be treated as being “uniformly aligned.” For example, if we adopt a 12.5 ms frame-shift to extract Mel spectrograms, considering the framerate of XLSR is 20ms, the length ratio of $R:MEL:Speech$ would be $5:8:1600$. In this work, we adopt an up-sampling factor of 8 and then down-sample by a factor of 5 to map discrete representations to Mel ones. Transposed convolution is used for up-sampling, followed by a Multi-Receptive Field Fusion (MRF) module, similar to the HiFi-GAN generator. With an up-sampling factor of 8, the configuration of the generator was changed for up-sampling rates to a sequence of (4, 2) with corresponding kernel sizes (12,8), while the hyper-parameters of residual blocks are the same as those in HiFi-GAN V1. For the down-sampling, we use average pooling. The decoder part is the same as that of txt2vec, which is an FFT. The decoder takes the output of average pooling and language and speaker embeddings to produce the Mel spectrograms. The optimization objective of this vec2mel is defined as: $\tiny\mathcal{L}_{mel}=\tiny{{\left\|\hat{MEL}-MEL\right\|}_{2}^{2}}$.
@@
The details of the architecture and operation of the multi-stage encoder are provided in [54, 55]. Here, we illustrate the audio waveform generation process with a 2-stage encoder as an example as shown in Fig. 4, which consists of the following steps: 1) The first-stage encoder receives discrete sequences $R$ and produces the first hidden representation sequences. These sequences are then downsampled and passed to the second-stage encoder to obtain the second hidden representation, which has a lower time resolution. 2) The second hidden representation sequences are quantized using codebook 2 to obtain $z^{(2)}$. These sequences are then up-sampled and combined with the first hidden representation sequences to obtain $z^{(1)}$ using codebook 1. 3) The residual output of $z^{(1)}$ and $z^{(2)}$ is used to reconstruct the waveform using the frame decoder and HiFi-GAN-based waveform generator.
@@ IV-A Dataset
To ensure the balance of language, speakers, and gender, we screened the aforementioned databases and combined them into our multilingual and multispeaker database as shown in Table I. In addition to the two multispeaker databases MLS and GLB, for which most of the speakers have about 100 audio samples, we also selected 100 audio samples for each language from three single-speaker databases CSS10, LJS, and NST. The reason that we also selected some data from a single-speaker database is to aim to make a fairer comparison with single-speaker and single-language models in the future, but this is not included in this work.
To eliminate the potential impact of varying sampling rates as a confounding variable, we resampled all audio to a consistent 16 kHz rate and applied amplitude normalization using sv56 [62]. We set the frame and hop sizes to 1,024 and 200 (12.5 ms), respectively, to extract the Mel spectrograms from 16 kHz raw speech.
TABLE I: Details of the training corpus. ##\##Spk represents the total number of speakers per gender and language. ##\##Sent represents the total number of utterances per gender and language. “MLS”, “GLB” and “Other” represent a number of speakers selected from each of the MLS dataset, the Global Phone dataset and three single-speaker datasets, respectively.
@@
en Female 46 12.06 3,521 45 00 1(LJ)
Male 45 11.87 3,420 45 00 0
fr Female 45 12.29 4,661 00 45 0
Male 46 11.88 4,775 00 45 1(CSS10)
ge Female 46 11.92 3,892 38 07 1(CSS10)
Male 45 12.47 6,576 00 45 0
pt Female 46 09.15 3,596 03 43 0
Male 45 09.76 3,793 04 41 0
@@
sw Female 45 09.57 4,866 00 45 0
Male 46 09.32 4,776 00 45 1(NST)
TABLE II: Details of the pre-trained model.
@@ IV-B Pre-trained Models
In addition to utilizing the above six languages’ paired datasets for direct training of our TTS system, to further incorporate multi-lingual knowledge into our system, we also employed two pre-trained multilingual models, XLSR-53222https://huggingface.co/facebook/wav2vec2-large-xlsr-53 and XPhoneBERT [^3] https://github.com/VinAIResearch/XPhoneBERT, trained on audio-only and text-only data respectively, as shown in Table II. For the speaker encoder, we use the publicly available ECAPA-TDNN model [^4] https://huggingface.co/speechbrain/spkrec-ecapa-voxceleb, which was trained with the multilingual VoxCeleb 1&2 dataset.
TABLE III: Experimental Setup. txt2intfeat and intfeat2wav represent text to intermediate features and intermediate features to waveform, respectively. SCL is speaker consistency loss. ECAPA and H/ASP are two different speaker encoders. Mel and SSL respectively represent the use of Mel spectrograms or self-supervised learning representations as intermediate features.
@@ IV-C Experimental Setup and Model Architecture
TABLE IV: MOS, DMOS, and SECS results for seen speakers among different systems. The gray color scale indicates the relative value, and darker color indicates a better result. MOS and DMOS are reported with 95% confidence intervals.
@@
TABLE V: MOS, DMOS, and SECS results for unseen speakers among different systems. The gray color scale indicates the relative value, and darker color indicates a better result. MOS and DMOS are reported with 95% confidence intervals.
@@
In summary, the differences of the various TTS systems are listed in Table III. Furthermore, for ZMM-TTS and FSM, we also compare different input representations, using different suffixes to indicate them. For example, ZMM-TTS1c, ZMM-TTS1i and ZMM-TTS1x indicate that the inputs to the ZMM-TTS1 model are characters, IPA, and pre-trained phoneme representations, respectively. The implementation details are summarized as follows and each module is trained independently:
  • Hifi-GAN: As for waveform generation in FSM1 and ZMM-TTS1, HiFi-GAN [^5] https://github.com/jik876/hifi-gan is chosen as our neural vocoder. This HiFi-GAN is trained for 2.5M steps on a single NVIDIA A100 GPU with a batch size of 16 using our 6 languages training data.
  • FastSpeech: Similar to text2vec in Section III-A, we also train a baseline system like FastSpeech as an acoustic model. This FastSpeech has the same encoder, decoder speaker, and language representations as text2vec, but the output is Mel representations rather than SSL ones. FastSpeech, text2vec, and vec2mel were all trained for 1.2M steps on a single NVIDIA A100 GPU with a batch size of 16.
@@
  • YourTTS: YourTTS is a multilingual system based on VITS which has achieved state-of-the-art (SOTA) results in zero-shot multispeaker TTS. The original YourTTS implementation uses the H/ASP model [63] as a speaker encoder. Therefore, we also train a YourTTSE, which means YourTTS using the same speaker encoder ECAPA as our ZMM-TTS model. Furthermore, there is a speaker consistency loss (SCL) in YourTTS to improve speaker similarity. Therefore, we also train a YourTTSW, in which “W” means YourTTS without SCL, as in our proposed model. In YourTTSW, ECAPA is also used as speaker encoder rather than H/ASP. All these YourTTS-based models were trained for 1,000k steps on a single NVIDIA A100 GPU with a batch size of 32, and we always select the best checkpoint based on the development set.

### v2

@@ III METHOD
Our proposed model is a multilingual and multispeaker synthesis system that uses self-supervised discrete speech representations from a pre-trained multilingual wav2vec 2.0 (XLSR-53) model. Quantization in XLSR-53 is based on product quantization by choosing quantized representations from codebooks $\bm{C}\in\mathbb{R}^{G\times M\times D}$ with $M=320$ entries each, $G=2$ are two separate codebooks in XLSR-53 and each quantization representation has $D=384$ dimensions.
Given a dataset $\mathcal{S}=\{x_{i},y_{i}\}$, where $y$ is an audio sample and $X=\{x_{0},x_{1},...,x_{T_{m}}\}$ is its corresponding text sequences, we use a pre-trained XLSR-53 model to encode each audio sample into discrete code index sequences $V=\{{v_{1}^{1:G},v_{2}^{1:G},...,v_{T_{c}}^{1:G}\}}$ and the corresponding discrete representation is $R=\{r_{1}^{1:G},r_{2}^{1:G},...,r_{T_{c}}^{1:G}\}$. Obviously, there is a one-to-one correspondence between $V$ and $R$, and once we obtain $V$, we can obtain $R$ by looking up codebooks $\bm{C}$ when $r=C[:,v,:]$.
Figure 1 depicts the architectures in which our proposed model can be implemented. It consists of two steps, as shown in the following formula. First, the discrete code index sequences $V$ are predicted from the text $x$ and the codebook of XLSR-53 $\bm{C}$ is looked up to obtain the discrete representations $R$, and then the waveform is predicted from the discrete representations $R$.
@@
where $S$ denotes the speaker representation and $L$ represents the language identity. Note that we discard the language input and related embedding layers, enabling direct inference on unseen languages.
The txt2vec module is similar to the acoustic model in prior speech synthesis approaches, so we constructed it on the basis of the popular acoustic model FastSpeech. Vec2wav is similar to a vocoder. To obtain a broader understanding of the potential complementarity and differences between SSL representations and Mel features, we proposed two distinct approaches to implement vec2wav. The first is to convert discrete representations $R$ into Mel spectrograms and then convert them through an additional vocoder, and the second is to directly convert $R$ into audio. Both methods make use of the up-sampling operation based on transposed convolution in HiFi-GAN [68], as in [54]. It should be noted that although there are $G$ different codebooks, in the vec2wav model, the input $R$ is represented by flattened representations from $G$ codebooks together. In the following two subsections, we will provide an explanation of the two components, respectively.
Figure 1: Overview of ZMM-TTS. The modules txt2vec and vec2wav are trained independently. Language ID is utilized for high-resource languages and few-shot adaptation, but it is not used for direct inference without fine-tuning.
@@ III-A Predicting Discrete Code Index from Text
Figure 2: Architecture of our proposed txt2vec model.
The txt2vec model is based on FastSpeech [39], a non-autoregressive (NAR) acoustic model with explicit duration modeling. In our work, we use a learnable aligner [69] instead of relying on the predicted phonemes durations of an autoregressive teacher TTS model. This aligner is more robust for long sequences. The decoder part is the same as that of FastSpeech, which is a Feed-Forward Transformer (FFT). For the encoder part, depending on the different types of input representations, it can be either FFT-like or BERT-like. Figure 2 shows the structure of txt2vec.
@@ III-A1 Input representations and text encoder
We attempt to use different input representations including characters, IPA, and pre-trained phoneme representations for multilingual TTS.
@@
III-A2 Multispeaker and multilingual control
@@ III-A3 Loss Function
In this txt2vec stage, the text encoder operates on the input tokens $x$ and produces hidden states, which, combined with language and speaker embeddings, are used to predict the duration $d$ by the duration predictor. The up-sampled encoder outputs, together with the language and speaker embeddings, are used as the decoder input. The decoder then generates discrete values for both $V$ and $R$.
@@
$\begin{split}\hat{r}_{t}^{g}=\sum_{i=1}^{M}p(v_{t}^{g}=i|h_{t}^{g})C_{g,i}\end{split}$ (4)
Eq. (4) is a weighted sum of the code vectors, where the weights are the predicted probabilities of choosing the codes. In Eq. (4), $p(v_{t}^{g}=i|h_{t}^{g})$ represents the probability of predicting the code index value $v_{t}^{g}$ as $i$ at time $t$. Here, $g\in G$ denotes the g-th codebook and $C_{g,i}$ denotes the vector corresponding to the i-th index in the g-th codebook at $C^{G\times M\times D}$. The regression L2 loss $\mathcal{L}_{MSE}$ between the predicted and targeted $R$ could be represented as
$\mathcal{L}_{mse}=\frac{1}{T}\sum_{t=1}^{T}\frac{1}{G}\sum_{g=1}^{G}{\left\|\hat{r}_{t}^{g}-r_{t}^{g}\right\|}^{2}_{2}$ (5)
@@
Note that since $p(v_{t}^{g})$ is a one-hot vector, the $r_{t}^{g}$ is only one entry in the codebook instead of being multiplied by softmax weights as in Eq. (5). Therefore, in inference, we directly select the maximum probability index in the $P(i|h_{t}^{g})$, and look up the corresponding $(\hat{r}_{t}^{g})$ in the code $C^{G\times M\times D}$ as the input of the next-stage model:
@@
III-B Predicting a waveform from discrete representations
Although most methods convert features directly into audio [54], spectral features may still have a few advantages in several low-resource tasks [70]. To obtain a broader understanding of the potential complementarity and differences between SSL representations and Mel features, we adopted two different ways to implement the vec2wav model: with and without an independent Mel-based vocoder as in the following formula,
@@ III-B Predicting a waveform from discrete representations
This vec2wav model consists of four parts: up-sampling, down-sampling, decoder, and pre-trained Mel-based vocoder as shown in Figure 3. To learn Mel spectrograms from $R$, the duration model is no longer necessary. Although the sequences $R$ and $Mel$ have different “sampling rates,” they can be treated as being “uniformly aligned.” For example, if we adopt a 12.5 ms frame-shift to extract Mel spectrograms, considering the framerate of XLSR is 20ms, the length ratio of ${\color[rgb]{0,0,0}R:Mel:Speech}$ would be $5:8:1600$. In this work, we adopt an up-sampling factor of 8 and then down-sample by a factor of 5 to map discrete representations to Mel ones. Transposed convolution is used for up-sampling, followed by a Multi-Receptive Field Fusion (MRF) module, similar to the HiFi-GAN generator. With an up-sampling factor of 8, the configuration of the generator was changed for up-sampling rates to a sequence of (4, 2) with corresponding kernel sizes (12,8), while the hyper-parameters of residual blocks are the same as those in HiFi-GAN V1. For the down-sampling, we use average pooling. The decoder part is the same as that of txt2vec, which is an FFT. The decoder takes the output of average pooling and language and speaker embeddings to produce the Mel spectrograms. The optimization objective of this vec2mel is defined as: ${\mathcal{L}_{Mel}={\left\|\hat{Mel}-Mel\right\|}_{2}^{2}}$.
@@
The details of the architecture and operation of the multi-stage encoder are provided in [71, 72]. Here, we illustrate the audio waveform generation process with a 2-stage encoder as an example as shown in Figure 4, which consists of the following steps: 1) The first-stage encoder receives discrete sequences $R$ and produces the first hidden representation sequences. These sequences are then downsampled and passed to the second-stage encoder to obtain the second hidden representation, which has a lower time resolution. 2) The second hidden representation sequences are quantized using codebook 2 to obtain $z^{(2)}$. These sequences are then up-sampled and combined with the first hidden representation sequences to obtain $z^{(1)}$ using codebook 1. 3) The residual output of $z^{(1)}$ and $z^{(2)}$ is used to reconstruct the waveform using the frame decoder and HiFi-GAN-based waveform generator.
@@ IV-A Dataset
To ensure the balance of language, speakers, and gender, we screened the aforementioned databases and combined them into our multilingual and multispeaker database as shown in Table II. In addition to the two multispeaker databases MLS and GLB, for which most of the speakers have about 100 audio samples, we also selected 100 audio samples for each language from three single-speaker databases CSS10, LJS, and NST. The reason that we also selected some data from a single-speaker database is to aim to make a fairer comparison with single-speaker and single-language models in the future, but this is not included in this work.
To eliminate the potential impact of varying sampling rates as a confounding variable, we resampled all audio to a consistent 16 kHz rate and applied amplitude normalization using sv56 [79]. We set the frame and hop sizes to 1,024 and 200 (12.5 ms), respectively, to extract the Mel spectrograms from 16-kHz raw speech.
TABLE II: Details of the training corpus. ##\##Spk represents the total number of speakers per gender and language. ##\##Sent represents the total number of utterances per gender and language. “MLS”, “GLB” and “Other” represent a number of speakers selected from each of the MLS dataset, the Global Phone dataset and three single-speaker datasets, respectively.
@@
en Female 46 12.06 3,521 45 0 1(LJ)
Male 45 11.87 3,420 45 0 0
fr Female 45 12.29 4,661 0 45 0
Male 46 11.88 4,775 0 45 1(CSS10)
ge Female 46 11.92 3,892 38 7 1(CSS10)
Male 45 12.47 6,576 0 45 0
pt Female 46 9.15 3,596 3 43 0
Male 45 9.76 3,793 4 41 0
@@
sw Female 45 9.57 4,866 0 45 0
Male 46 9.32 4,776 0 45 1(NST)
TABLE III: Details of the pre-trained model.
@@ IV-B Pre-trained Models
In addition to utilizing the aforementioned six languages’ paired datasets to train our TTS system directly, to incorporate more multilingual knowledge into our system, we also employed two pre-trained multilingual models, XLSR-53222https://huggingface.co/facebook/wav2vec2-large-xlsr-53 and XPhoneBERT [^3] https://github.com/VinAIResearch/XPhoneBERT, trained on audio-only and text-only data respectively, as shown in Table LABEL:tab:pretrained. For the speaker encoder, we use the publicly available ECAPA-TDNN model [^4] https://huggingface.co/speechbrain/spkrec-ecapa-voxceleb, which was trained with the multilingual VoxCeleb 1&2 dataset.
TABLE IV: Experimental Setup. txt2intfeat and intfeat2wav represent text to intermediate features and intermediate features to waveform, respectively. SCL is speaker consistency loss. ECAPA and H/ASP are two different speaker encoders. Mel and SSL respectively represent the use of Mel spectrograms or self-supervised learning representations as intermediate features.
@@ IV-C Experimental Setup and Model Architecture
TABLE V: MOS, DMOS, and SECS results for seen speakers among different systems. The gray color scale indicates the relative value, and darker color indicates a better result. MOS and DMOS are reported with 95% confidence intervals.
@@
TABLE VI: MOS, DMOS, and SECS results for unseen speakers among different systems. The gray color scale indicates the relative value, and darker color indicates a better result. MOS and DMOS are reported with 95% confidence intervals.
@@
In summary, the differences of the various TTS systems are listed in Table IV. Furthermore, for ZMM-TTS and FSM, we also compare different input representations, using different suffixes to indicate them. For example, ZMM-TTS1c, ZMM-TTS1i and ZMM-TTS1x indicate that the inputs to the ZMM-TTS1 model are characters, IPA, and pre-trained phoneme representations, respectively. The implementation details are summarized as follows and each module is trained independently:
  • HiFi-GAN: As for waveform generation in FSM1 and ZMM-TTS1, HiFi-GAN [^5] https://github.com/jik876/hifi-gan is chosen as our neural vocoder. This HiFi-GAN is trained for 2.5M steps on a single NVIDIA A100 GPU with a batch size of 16 using our six languages training data.
  • FastSpeech: Similar to txt2vec in Section III-A, we also train a baseline system like FastSpeech as an acoustic model. This FastSpeech has the same encoder, decoder speaker, and language representations as txt2vec, but the output is Mel representations rather than SSL ones. FastSpeech, txt2vec, and vec2mel were all trained for 1.2M steps on a single NVIDIA A100 GPU with a batch size of 16. Both FastSpeech and txt2vec were built from an open-source repository [^6] https://github.com/keonlee9420/Comprehensive-Transformer-TTS.
@@
  • YourTTS: YourTTS is a multilingual system based on VITS that has achieved state-of-the-art (SOTA) results in zero-shot multispeaker TTS. The original YourTTS implementation uses the H/ASP model [80] as a speaker encoder. Therefore, we also train a YourTTSE, which means YourTTS using the same speaker encoder ECAPA as our ZMM-TTS model. Furthermore, there is a speaker consistency loss (SCL) in YourTTS to improve speaker similarity. Therefore, we also train a YourTTSW, in which “W” means YourTTS without SCL, as in our proposed model. In YourTTSW, ECAPA is also used as a speaker encoder rather than H/ASP. All these YourTTS-based models were trained for one million steps on a single NVIDIA A100 GPU with a batch size of 32, and we always select the best checkpoint based on the development set.